Fermi National Laboratory
 

Accelerator Update for March 19, 2007 - March 21, 2007

 

Monday March 19

The day shift began with Operations monitoring the Tevatron store 5290 (86.04E30) and the Recycler stash (178.5E10), with the Antiproton Source (Pbar) stacking (7.0E10), with the Booster, Linac, Main Injector (MI), and Recycler operating normally, with MiniBooNE off, with NuMI and the Meson MTest experiment taking beam, and with D-Zero (D0) and the Collider Detector at Fermilab (CDF) using the colliding beams.

Pbar experts began stacktail equalizer studies at 8:12 AM.

A Booster RF station (BRF6) tripped off. It reset.

At 11:12 AM, Operations reported that the lithium lens would not turn on due to an external interlock. Operators reset this locally. Also, an Accumulator RF station (ARF1) wouldn’t ramp. Operators investigated and found that a low-level RF switch was off (A:R1LLEM). They turned it on and that allowed ARF1 to ramp.

At 3:03 PM, the Booster West Anode Power Supply (WAPS) tripped off. It reset.

At 3:10 PM, Operations reported that BRF17 & 18 had tripped off. Operators got the stations back on, but BRF18 was sparking a lot.

At 3:13 PM, WAPS tripped off again due to BRF18. RF experts bypassed the station.

Pbar studies ended at 3:36 PM.

At 4:17 PM, operators had to reboot the BPMs for the Pbar AP1 transfer line.

At 4:53 PM, Operations reported that the AP1 BPM VME wasn’t communicating properly. Operators couldn’t reset the VME, so they called in experts.

At 5:10 PM, Operations reported that the Preaccelerator 750keV buncher tripped off. It reset.

The MTest experimenters reported at 6:32 PM that they were done with the beam for the night.

Experts reported at 7:07 PM that they had resolved the AP1 BPM communications problem.

Operations reported at midnight that the Recycler had stashed antiprotons twice during the evening shift.

Tuesday March 20

The midnight shift began with Operations monitoring store 5290 (40.27E30) and the stash (277.0E10), with Pbar stacking (38.8E10), and with NuMI taking beam.

Operations terminated store 5290 at 5:02 AM.

A TeV expert began a TeV study at 6:24 AM.

MTest resumed taking beam at 6:43 AM.

Operation put the TeV into shot setup at 7:29 AM.

Operations reported at 7:59 AM that the Recycler had stashed three times during the midnight shift.

At 8:23 AM, Operations reported a possible Pbar AP3 transfer line LCW leak.

Operations established store 5292 at 9:25 AM with an initial luminosity of 128.57E30. Some antiprotons were lost during the shot, making this a 36x32 store (The first number is for bunches of protons and the second number is for antiprotons). Later, TeV experts reported that their antiproton closure program had a problem, which was another reason why the luminosity was so low. They continued their investigation.

Operations prepared Booster at 10:22 AM for an access to investigate BRF18 problems.

Operations had MI-65 prepared for an access at 10:34 AM.

Cryo system technicians began work on the Pbar dry engine at 10:35 AM.

Booster RF experts reported at 11:06 AM that BRF18 problems would take a couple of hours to repair, which would be done at a later date. Also, Pbar experts reported that the AP3 LCW leak on IQ19 was repaired.

Cryo system techs reported at 11:11 AM that they were completing their repairs on the Pbar dry engine.

Booster came out of access at 11:15 AM.

MI-65 came out of access at 11:24 AM.

Pbar came out of access at 11:38 AM.

Pbar, NuMI, and MTest resumed taking beam by 11:55 AM.

At 2:31 PM, a vacuum expert reported that the TeV sector B23 vacuum pump was failing and needed to be replaced.

Operations reported at 3:56 PM that the Recycler had stashed three times during the day shift.

A Recycler Electron Cooling expert reported at 4:25 PM that the Pelletron would be off for 2 -3 hours for LCW work.

Operations reported that beam to MTest was turned off at 6:05 PM.

An expert reported at 7:40 PM that the Pelletron was fully operational; they had cleaned a clogged LCW strainer.

Pbar experts began a target lens gradient study at 7:57 PM.

At 9:46 PM, a Linac RF station (LRF3) tripped off for the first of three times in 44 minutes. It reset.

Operations reported at midnight that the Recycler had stashed once during the evening shift.

Wednesday March 21

The midnight shift began with Operations monitoring store 5292 (42.09E30) and the stash (174.1E10), with Pbar stacking (96.0E10), and with NuMI taking beam.

Pbar experts completed their study at 12:59 AM.

At 5:18 AM, BRF7 & 8 tripped off due to a vacuum pump (IPRF17) trip. Everything reset okay.

At 5:22 AM, Operations reported problems with the Recycler transverse emittances. Experts helped tune the beam.

Operations terminated store 5292 at 7:29 AM.

Operations put the TeV into shot setup at 7:43 AM.

Operations reported at 8:05 AM that the Recycler had stashed three times during the midnight shift.

The Plans for Wednesday

The plan for today is to establish a new store this morning. The Run Coordinator will allow Recycler, Pbar, and MI studies.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 03/21/2007    email Fermilab